• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Moffit

    Log entry time 22:19:28 on June15,2004

    Entry number 125899

    keyword=New BCM7 and BCM8 calibration run #2171

    Used run 2171 as a calibration run for bcm7 and bcm8 (upstream and downstream x3 BCMs from the 10kHz downconverter).  Rough BCM (as stated in EPICS hac_bcm_average) values are recorded in scandata1.
    
    Here are some rough event numbers of what happened during the run:
    
    Ev Range       Current
    -----------------------
       0k- 6.6k       40
     6.8k-12.0k        0
    17.0k-23.0k       30
    25.0k-30.0k        0
    35.0k-41.0k       20
    41.3k-47.7k        0
    50.0k-56.0k       10
    56.5k-62.5k        0
    70.6k-END          5
    
    This data was taken with raster on (3.5x3.6) on the Carbon target.  BLumi HVs were adjusted in run 2170 s.t. they were high enough in the ADCs during the entire run 2171.  Left Septum magnet was ramping during most of this run.